Three particles lie on the x axis: particle 1, with a charge of 1 x 10–8 C is at x = 1 cm, particle 2, with a charge of 2 x 10–8 C, is at x = 2 cm, and particle 3, with a charge of -3 x 10–8 C, is at x = 3 cm. The potential energy of this arrangement, relative to the potential energy for infinite separation, is:
Answer #1

Solutions) V 03 + o i as 2 3 Y-> het q se=1 cm or reg 2cm = 1810-8c = 2x10-8c - 3X 1080 2 300 Org *3 il ) Patential evergy fo4 VE 4.95X10 J value This is required the 아 giren of the potential energy configuration.